OpenStudy (anonymous):

Greatest Common Factor & Least Common Multiple 12 a^5 b and 40 a^5 b^5

OpenStudy (anonymous):

GCF: Work out the numbers first. GCF of 12 and 40 is 2^2, which is 4. Then, work out the rest. GCF of a^5 * b and a^5 * b^5 is a^5 * b. So, the GCF is 4 * a^5 * b LCM: Work out the numbers first. LCM of 12 and 40 is 2^3 * 3 * 5, which is 120. Then, work out the rest. LCM of a^5 * b and a^5 * b^5 is a^5 * b^5. So, the LCM is 120 * a^5 * b^5.

OpenStudy (anonymous):

i have one more can you help me with

OpenStudy (anonymous):

15 f^2 g h^3 and 60 f^2 h^3

OpenStudy (anonymous):

You should take the same approach. GCF GCF of 15 (=3*5) and 60 (=2^2*3*5) is 15, and GCF of f^2*g*h^3 and f^2*h^3 is f^2*h^3. So, GCF is 15*f^2*h^3. LCM LCM of 15 (=3*5) and 60 (=2^2*3*5) is 60, and LCM of f^2*g*h^3 and f^2*h^3 is f^2*g*h^3. So, LCM is 60*f^2*g*h^3. I hope this helps.